OpenStudy (anonymous):

determine (f*h)(3) when f(x) = {(3,5)(2,4)(1,7)} h(x)= {(3,2)(4,3)(1,6)}

OpenStudy (barrycarter):

(f*h)(3) = f(3)*h(3) = 5*2 = 10